Find the LCD and add the fractions.
1/125 + 4/50 =

Found 2 solutions by lynnlo, Alan3354:
Answer by lynnlo(4176)
Answer by Alan3354(69443)

Find the LCD and add the fractions.
1/125 + 4/50
---------
LCD = 250
----
1/125 + 4/50 = (2 + 20)/250 = 22/250
= 11/125